OpenStudy (anonymous):

what are the solutions of the equation l 2-3x l=8x ? i need help, i don't know how to solve for absolute value problems

hero (hero):

-2+3x = 8x 2-3x = 8x

hero (hero):

Solve both for x

OpenStudy (anonymous):

the thing is, i was given the solutions, but i get 2/11 but thats not one of the answers listed

OpenStudy (anonymous):

yeah i got that, but its not listed as one of the possible answers :/

hero (hero):

Well, I can't help you with that. I can only present to you the answer based on algebra. If your book presents something else, I can't do anything about that.

hero (hero):

Oh wait...

hero (hero):

Try this: \[(2-3x)^2 = (8x)^2\]

hero (hero):

That's the alternative way of approaching this

hero (hero):

Weird thing is, even if you did that, you still get the same answer, so I don't know

hero (hero):

according to wolfram alpha, 2/11 is the only possibility for absolute value. If I were you, I would consult your teacher regarding this.

hero (hero):

The only other way to present the answer is in decimal form

hero (hero):

.18 repeating